OpenStudy (anonymous):

will medal and fan!! Which of these is a simplified form of the equation 9p + 3 = − p + 7 + 2p + 3p?

OpenStudy (anonymous):

Where are the options?

OpenStudy (anonymous):

Which of these is a simplified form of the equation 9p + 3 = − p + 7 + 2p + 3p? 5p = 4 1 5p = 10 9 = 4 12 = 13

OpenStudy (anonymous):

Add and subtract like terms and you end up with: \[9p+3=4p+7\] Subtract 4p from both sides and you end up with: \[5p+3=7\] Subtract 3 from both sides and you end up with: \[5p=4\] That's the answer!